OpenStudy (dls):

a boy travels 10m due north and then 7m due east.What is his displacement+angle? I got displacement as 12.21 metres,how do i get the angle?

OpenStudy (dls):

i did tan theta=10/7=1.42 but maybe its wrong..

OpenStudy (anonymous):

arctan(Y-axis/X-axis) = arctan(10/7) = 55.01

OpenStudy (anonymous):

the angle is correct if taken north of east.